Please help me with this question "let the width the length be real numbers. The perimeter remains a constant 24 cm. Draw the graph for 1_
Answer by Theo(13342)

perimeter is equal to 2 times the length plus 2 times the width for a rectangle.

let L equal the length and W equal the width.

your equation is 2L + 2W = 24

divide both sides of this equation by 2 to get L + W = 12

solve for L to get L = 12 - W.

your domain is 1 <= W <= 11.

if you let y equal L and x equal W, then your graph will look like this:

$$$

when x = 1, then y = 11, and 2x + 2y = 24

when x = 2, then y = 10, and 2x + 2y = 24

when x = 3, then y = 9, and 2x + 2y = 24

in fact, all combinations of x and the corresponding value of y in the equation of y = 12 - x will yield 24 as the answer when you place them in the equation of 2x + 2y = z, where z represents the perimeter of the rectangle.